# Disclosure Avoidance for the 2020 Census Demographic and Housing Characteristics File

### v1

@@ 4.1. DHCH Production PLB Allocations
    HHTYPE_RECODE. 183 levels: Defined as the cross product of HHTYPE attributes, MULTIGEN $\times$×SPOUSE×$\times$×SIZE×$\times$×OWN_CHILD.
@@
Query ID
Marginal Query Group Name
Cells
1H
MULTIGEN×$\times$×HISPANIC×$\times$×TENURE_2LEVELS
8
2H
HISPANIC×$\times$×TENURE_2LEVELS
4
3H
HISPANIC×$\times$×TENURE_2LEVELS×$\times$×RACE
28
4H
PARTNER_TYPE_CHILD_STAT×$\times$×SEX×$\times$×TENURE_2LEVELS
48
5H
SPOUSE×$\times$×HISPANIC×$\times$×TENURE_2LEVELS
20
6H
SEX×$\times$×HISPANIC×$\times$×TENURE×$\times$×RACE×$\times$×HHTYPE_RECODE
15,372
7H
SEX×$\times$×HISPANIC×$\times$×TENURE×$\times$×RACE×$\times$×HHAGE×$\times$×HHTYPE_RECODE
1,383,48
8H
TENURE×$\times$×HHAGE×$\times$×HHTYPE_RECODE
4,941
9H
SEX×$\times$×HISPANIC×$\times$×TENURE×$\times$×RACE×$\times$×ELDERLY×$\times$×HHAGE×$\times$×HHTYPE
1,578,528
10H
TENVACGQ
35
Each marginal query group in Table 2 is assigned a query ID, and the PLB allocations for the production US DHCH implementation for each query ID and each geographic level are given in Table 3. The values in the table below are the actual $\rho$ allocations multiplied by 10,000; in other words, the actual $\rho$ allocations used in the DHCH DAS production settings are given by the values in the table below divided by 10,000. The total value of $\rho$ allocated to all marginal query groups for our production DHCH settings was 15401/2000 ($\approx 7.70$).
Table 3. Production US DHCH $\rho$ Allocations ×10,000$\times$ 10,000× 10 , 000 (i.e., divide by 10,000 to obtain the $\rho$ value)
Query ID
US
State
County
Prim
TSG
TS
OBG
Block
1H
792
714
986
973
973
1,150
291
11
2H
212
193
266
214
214
310
78
3
3H
212
191
1,500
1,500
1,500
1,750
442
17
4H
792
1428
1,972
1,946
1,946
2,300
581
22
5H
792
714
3,944
973
973
4,601
1,162
44
6H
257
3,216
1,380
973
973
1,610
407
15
7H
257
3,216
1,380
973
973
1,610
407
15
8H
212
193
266
214
214
310
78
3
9H
1,049
3,931
1,380
973
973
1,610
407
15
10H
259
3,216
1,380
973
973
1,610
407
15
@@ 4.2. DHCP Production PLB Allocations
As described in the previous section for the DHCH schema, the marginal query groups used in the DHCP strategies are provided in Table 4. Many of these marginals are defined by the attributes in the Section 2.2. The definitions of the remaining marginals, i.e., recodes, are given in the list below.
@@
    AGE_38_GROUPS. 38 levels: Indicates which of the following categories the respondent’s age is in AGE $=0,$ AGE $=1,\dots,$ AGE $=14,$ AGE $\in[15,17],$ AGE $\in[18,19],$ AGE $\in[20,24],$ AGE $\in[25,29],\dots,$ AGE $\in[55,59],$ AGE $\in[60,61],$ AGE $\in[62,64],$ AGE $\in[65,66],$ AGE $\in[67,69],$ AGE $\in[70,74],$ AGE $\in[75,79],\dots,$ or AGE $\geq 110.$
@@
    GQ_CONSTR_GROUPS. 6 levels: Indicates if the respondent is a resident of a GQ with a lower-bound constraint on AGE given by 0, 3, 15, 16, 17 years of age, as described in Table 1, or if the respondent is a member of a household. In other words, GQ_CONSTR_GROUPS indicates if the respondent’s GQ type is in $\{201,202,203,301,401,402,403,701\},\{405\},$$\{101,102,103,104,105\},$$\{501,801,802,900,901,702,704,706,903,904\},\{106,404,601,602\},$ or if the respondent is in a household.
@@
    AGE_10_GROUPS. 10 levels: Indicates which of the following categories the respondent’s age is in AGE $\leq 3,$ AGE $=[3,14],$ AGE $=15,$ AGE $=16,$ AGE $=17,$ AGE $=[18,19],$ AGE $\in[20,24],$ AGE $\in[25,29],$ AGE $\in[30,34],$ or AGE $\geq 35.$
@@
    AGE_40_GROUPS. 40 levels: Indicates which of the following categories the respondent’s age is in AGE $=0,$ AGE $=1,\dots,$ AGE $=17,$ AGE $\in[18,19],$ AGE $\in[20,24],$ AGE $\in[25,29],\dots,$ AGE $\in[55,59],$ AGE $\in[60,61],$ AGE $\in[62,64],$ AGE $\in[65,66],$ AGE $\in[67,69],$ AGE $\in[70,74],$ AGE $\in[75,79],\dots,$ AGE $\in[105,109],$ or AGE $\geq 110.$
@@
Query ID
Marginal Query Group Name
Cells
1P
AGE_18_64$\times$×RELGQ_4_GROUPS
12
2P
AGE_18_64$\times$×SEX
6
3P
AGE_38_GROUPS×$\times$×SEX
76
4P
HISPANIC×$\times$×SEX
4
5P
SEX×$\times$×RELGQ_4_GROUPS
8
6P
GQ_CONSTR_GROUPS×$\times$×AGE_10_GROUPS
60
7P
POP_SEHSD_TARGETS_RELSHIP
19
8P
HISPANIC×$\times$×SEX×$\times$×AGE_40_GROUPS×$\times$×RELSHIP_GQ×$\times$×CENRACE
252,000
9P
RELGQ×$\times$×AGE_40_GROUPS×$\times$×HISPANIC×$\times$×CENRACE×$\times$×SEX
423,360
10P
RELGQ×$\times$×SEX×$\times$×AGE×$\times$×HISPANIC×$\times$×CENRACE
1,227,744
@@
Table 5. Production US DHCP $\rho$ Allocations ×10,000$\times$ 10,000× 10 , 000
Query ID
US
State
County
Prim
TSG
TS
OBG
Block
1P
73
999
310
478
478
868
430
11
2P
73
999
310
478
478
868
430
11
3P
73
999
310
478
478
868
430
11
4P
73
999
310
478
478
868
430
11
5P
73
999
310
478
478
868
430
11
6P
292
3,996
1,240
1,912
1,912
3,472
430
11
7P
73
999
1,240
478
478
3,472
430
11
8P
73
999
310
478
478
868
430
11
9P
73
999
310
478
478
868
430
11
10P
73
999
310
478
478
868
430
11
@@ 5. Summary Metrics for 2020 DHC Production Implementations
[25] provides data users with $\rho-$zCDP summary metrics for the 2020 DHC release using a discrete Gaussian mechanism. The total PLB for all the summary metrics in this release was $\rho=0.7.$ We used a formally private method to produce these metrics because they summarize errors of the 2020 DHC release MDF-based tabulations with respect to CEF-based tabulations. This section provides the approximate mean absolute error (MAE) $\rho-$zCDP metrics from this document for several important DHC tabulations. The MAE of a query in a given geographic level is defined by first finding the query error in each geographic unit in the geographic level; afterward, the MAE is defined as the arithmetic mean of the absolute value of these errors. The PLB allocated to each of the state geographic level MAEs was set so that the reported approximate MAE was within 0.5 units of the unprotected CEF-based MAE value with probability 0.9. Likewise, the PLB allocated to the remaining summary metrics was set so that the reported approximate MAE was within 0.15 units of the unprotected CEF-based MAE value with probability 0.9. Many other $\rho-$zCDP approximate summary metrics in addition to the ones that are provided here are available in [25].
Tables 6 through 7 provide the MAE values of tabulations in the DHCH universe. As described previously, counts in this universe can be viewed as counts of households; alternatively, since there is one householder per household, counts in this universe can also be viewed as counts of householders. For the DHCP universe, Tables 8 and 9 provide the MAE of counts of individuals grouped by relationship to the householder and by their age group, respectively.
Table 6. Production US 2020 DHCH: Presence of Own Child of Householder MAE
Group
States
Counties
Places
Tracts
Own children under 6 years only
10.39
5.49
2.99
4.10
Own children between 6 and 17 years only
10.98
5.71
3.30
4.59
Own chilren under 6 and between 6 and 17 years
14.88
5.71
3.11
4.09
No own children under 18 years
12.84
6.05
4.03
5.20
Table 7. Production US 2020 DHCH: Tenure by Race MAE
Group
States
Counties
Places
Tracts
@@
White alone
22.78
5.56
2.70
3.27
Black/African American alone
27.76
4.34
1.54
2.30
American Indian/Alaska Native alone
28.88
3.40
1.01
1.50
Asian alone
19.39
3.33
1.19
1.99
Native Hawaiian/OPI* alone
15.86
1.75
0.38
0.47
Some Other Race alone
13.33
3.81
1.60
2.24
Two or More Races
15.08
4.20
2.12
2.74
@@
White alone
17.08
4.70
2.26
2.95
Black/African American alone
13.92
3.94
1.42
2.26
American Indian/Alaska Native alone
20.80
2.83
0.99
1.47
Asian alone
12.10
2.67
0.95
1.63
Native Hawaiian/OPI* alone
13.47
1.59
0.42
0.61
Some Other Race alone
11.82
3.29
1.41
2.15
Two or More Races
12.24
3.77
1.66
2.49
@@
Table 8. Production US 2020 DHCP: Relationship to Householder MAE
Group
States
Counties
Places
Tracts
Householder
23.02
6.70
4.32
4.26
Opposite-sex husband/wife/spouse
8.43
4.50
3.39
3.26
Opposite-sex unmarried partner
10.06
3.45
2.40
2.37
Same-sex husband/wife/spouse
8.39
2.58
1.44
1.47
Same-sex unmarried partner
6.92
2.33
1.37
1.35
Biological son or daughter
18.67
5.92
3.75
3.68
Adopted son or daughter
9.98
3.64
2.17
1.98
Stepson or stepdaughter
9.59
3.75
2.29
2.17
Brother or sister
8.35
3.30
2.15
2.18
Father or mother
7.02
3.08
2.08
2.06
Grandchild
11.35
3.91
2.60
2.58
Parent-in-law
6.25
2.63
1.47
1.57
Son-in-law or daughter-in-law
5.31
2.70
1.63
1.67
Other relative
11.02
3.93
2.33
2.33
Roommate or housemate
9.73
3.80
2.15
2.30
Foster child
6.61
2.58
1.46
1.23
Other nonrelative
10.80
3.84
2.37
2.21
Table 9. Production US 2020 DHCP: 5 Year Age Groups MAE
Group
Counties
Places
Tracts
AIR/ORT*
Under 5 years
9.52
4.47
5.28
3.95
5 to 9 years
10.64
4.36
5.84
4.35
10 to 14 years
10.63
4.50
5.95
4.61
15 to 19 years
8.31
4.52
4.39
2.84
20 to 24 years
4.57
2.78
3.06
2.08
25 to 29 years
4.59
2.80
3.09
2.17
30 to 34 years
4.39
2.77
3.02
2.12
35 to 39 years
6.56
3.01
4.06
3.29
40 to 44 years
6.36
2.94
4.00
2.86
45 to 49 years
6.25
2.93
3.95
2.89
50 to 54 years
6.22
2.99
3.96
3.20
55 to 59 years
6.17
3.07
3.96
3.35
60 to 64 years
8.08
3.80
4.99
3.33
65 to 69 years
7.90
3.63
4.75
3.53
70 to 74 years
6.01
2.84
3.67
2.83
75 to 79 years
5.87
2.71
3.50
2.76
80 to 84 years
5.88
2.55
3.33
2.54
85 to 89 years
5.39
2.29
3.10
1.90
90 to 94 years
5.29
2.19
2.75
1.55
95 to 99 years
4.77
1.69
2.00
0.99
100 to 104 years
3.14
0.94
0.77
0.38
105 to 109 years
1.29
0.27
0.14
0.40
over 109 years
0.64
0.03
0.05
0.19
@@ 6. Conclusion and Progress after the 2020 Census DHC Release
As described in Section 1, one important advantage of formally private mechanisms like the DAS is that they are sufficiently transparent to allow users to construct confidence intervals about the private CEF-based tabulations in a straightforward manner. To describe how this can be done, more background on how the strategy matrices described in the last section are used within the DAS is helpful. Specifically, after spine optimization, the noisy measurements for each tabulation and each geographic unit are defined by adding discrete Gaussian noise to the CEF-based tabulation of the given geographic unit. If the PLB allocation of the geographic unit and tabulation pair is given by $\rho,$ the mean-zero noise added to each level of the CEF-based tabulation is given by a realization of a random variable from the discrete Gaussian distribution, $N_{\mathbb{Z}}(\mu=0,\sigma^{2}=1/\rho),$ which is defined so that its probability mass function is proportional to $f(x)=\exp(-x^{2}/2\sigma^{2})$ for each $x\in\mathbb{Z}$ and zero otherwise [17]. We call the collection of all such noisy measurements for a given DAS implementation the noisy measurement files (NMFs). The NMFs provide sufficient information to construct the likelihood function for each geographic unit and each tabulation for which a noisy measurement is observed. This in turn can be used to generate point estimates and confidence intervals via simulation; alternatively, the underlying discrete Gaussian random variables can be modeled using continuous Gaussian random variables to derive closed form confidence intervals; see for example [20].
To support research and development on tools that would provide users inferences on underlying CEF-based tabulations, the Census Bureau has released NMFs for both the final 2010-CEF-based DDP production settings DAS implementations and the 2020 production DAS implementations as experimental data products. [26] provides instructions for downloading the NMFs of both of these implementations. Note that this method of producing confidence intervals using the NMFs typically does not result in an interval that is centered around the count implied by the MDF output from the DAS. Instead of relying on tabulations of counts from the MDF that is output from the DAS for the center of these confidence intervals, these intervals are centered around the inverse-variance-weighted mean of values derived from the NMFs, which, unlike MDF-based tabulations, provides an unbiased estimate of the count derived from the CEF.
The DAS team is also developing methods that provide confidence intervals for tabulations derived using the output of the production Redistricting and DHC MDFs. These confidence intervals are centered around the tabulations derived using the production MDFs, so they also provide a measure of the uncertainty due to disclosure avoidance of the published DHC tabulations.

### v2

@@ 5.1. DHCH Production PLB Allocations
[…]
@@
Query ID Marginal Query Group Name Cells
[…]
10H TENVACGQ 35
Each marginal query group in Table 3 is assigned a query ID, and the PLB allocations for the production US DHCH implementation for each query ID and each geographic level are given in Table 4. The values in the table below are the actual $\rho$ allocations multiplied by 10,000. The total value of $\rho$ allocated to all marginal query groups for our production DHCH settings was 15401/2000 ($\approx 7.70$).
Table 4. Production US DHCH $\rho$ Allocations $\times 10,000$ (i.e., divide by 10,000 to obtain the $\rho$ value)
Query ID US State County Prim* TSG* TS* OBG* Block
1H 792 714 986 973 973 1,150 291 11
2H 212 193 266 214 214 310 78 3
3H 212 191 1,500 1,500 1,500 1,750 442 17
4H 792 1428 1,972 1,946 1,946 2,300 581 22
5H 792 714 3,944 973 973 4,601 1,162 44
6H 257 3,216 1,380 973 973 1,610 407 15
7H 257 3,216 1,380 973 973 1,610 407 15
8H 212 193 266 214 214 310 78 3
9H 1,049 3,931 1,380 973 973 1,610 407 15
10H 259 3,216 1,380 973 973 1,610 407 15
@@ 5.2. DHCP Production PLB Allocations
As described in the previous section for the DHCH schema, the marginal query groups used in the DHCP strategies are provided in Table 5. Many of these marginals are defined by the attributes in Section 3.2. The definitions of the remaining marginals, i.e., recodes, are given in the list below.
@@
    AGE_38_GROUPS. 38 levels: Indicates which of the following categories the respondent’s age is in AGE =0,=0, AGE $1,\dots,$ AGE =14,=14, AGE $\in[15,17],$ AGE $\in[18,19],$ AGE $\in[20,24],$ AGE $\in[25,29],\dots,$ AGE $\in[55,59],$ AGE $\in[60,61],$ AGE $\in[62,64],$ AGE $\in[65,66],$ AGE $\in[67,69],$ AGE $\in[70,74],$ AGE $\in[75,79],\dots,$ or AGE $\geq 110.$
@@
    GQ_CONSTR_GROUPS. 6 levels: Indicates if the respondent is a resident of a GQ with a lower-bound constraint on AGE given by 0, 3, 15, 16, 17 years of age, as described in Table 2, or if the respondent is a member of a household. In other words, GQ_CONSTR_GROUPS indicates if the respondent’s GQ type is in $\{201,202,203,301,401,402,403,701\},\{405\},$ $\{101,102,103,104,105\},\{501,801,802,900,901,702,704,706,903,904\},\{106,404,601,$ 602},602\}, or if the respondent is in a household.
@@
    AGE_10_GROUPS. 10 levels: Indicates which of the following categories the respondent’s age is in AGE <3,<3, AGE =[3,14],=[3,14], AGE =15,=15, AGE =16,=16, AGE =17,=17, AGE =[18,19],=[18,19], AGE $\in[20,24],$ AGE $\in[25,29],$ AGE $\in[30,34],$ or AGE $\geq 35.$
@@
    AGE_40_GROUPS. 40 levels: Indicates which of the following categories the respondent’s age is in AGE =0,=0, AGE $1,\dots,$ AGE =17,=17, AGE $\in[18,19],$ AGE $\in[20,24],$ AGE $\in[25,29],\dots,$ AGE $\in[55,59],$ AGE $\in[60,61],$ AGE $\in[62,64],$ AGE $\in[65,66],$ AGE $\in[67,69],$ AGE $\in[70,74],$ AGE $\in[75,79],\dots,$ AGE $\in[105,109],$ or AGE $\geq 110.$
@@
Query ID Marginal Query Group Name Cells
[…]
7P POP_SEHSD_TARGETS_RELSHIP 19
[…]
@@
Table 6. Production US DHCP $\rho$ Allocations $\times 10,000$
Query ID US State County Prim* TSG* TS* OBG* Block
1P 73 999 310 478 478 868 430 11
2P 73 999 310 478 478 868 430 11
3P 73 999 310 478 478 868 430 11
4P 73 999 310 478 478 868 430 11
5P 73 999 310 478 478 868 430 11
6P 292 3,996 1,240 1,912 1,912 3,472 430 11
7P 73 999 1,240 478 478 3,472 430 11
8P 73 999 310 478 478 868 430 11
9P 73 999 310 478 478 868 430 11
10P 73 999 310 478 478 868 430 11
@@ 6. Summary Metrics for 2020 DHC Production Implementations
This section provides $\rho$-zCDP approximations of mean absolute error (MAE) metrics for several important DHC tabulations. The MAE of a query in a given geographic level is defined by first finding the query error in each geographic unit in the geographic level; afterward, the MAE is defined as the arithmetic mean of the absolute value of these errors. After computing the MAE values reported in this section, a discrete Gaussian mechanism was used to ensure the release satisfies $\rho$-zCDP. The PLB allocated to each of the state geographic level MAEs was set so that the reported approximate MAE was within 0.5 units of the unprotected CEF-based MAE value with probability 0.9. Likewise, the PLB allocations of the remaining summary metrics were set so that the reported approximate MAE was within 0.15 units of the unprotected CEF-based MAE value with probability 0.9. Many other $\rho$-zCDP approximate summary metrics, in addition to the ones that are provided here, are available in [DHCmetrics]. The total PLB for all the summary metrics in this release was $\rho=0.7.$
Table 7. Production US 2020 DHCH: Presence of Own Child of Householder $\rho$-zCDP MAE
Group States Counties Places Tracts
Own children under 6 years only 10.39 5.49 2.99 4.10
Own children between 6 and 17 years only 10.98 5.71 3.30 4.59
Own chilren under 6 and between 6 and 17 years 14.88 5.71 3.11 4.09
No own children under 18 years 12.84 6.05 4.03 5.20
Table 8. Production US 2020 DHCH: Householder Couple Type $\rho$-zCDP MAE
Group States Counties Places Tracts
Opposite-sex married couple household 10.67 3.43 3.01 2.89
Same-sex married couple household 8.33 2.50 1.47 1.59
Opposite-sex unmarried partner household 10.00 3.39 2.38 2.46
Same-sex unmarried partner households 8.41 2.37 1.28 1.43
Table 9. Production US 2020 DHCH: Householder Race $\rho$-zCDP MAE
Group States Counties Places Tracts
White alone 36.20 8.50 4.11 4.56
Black/African American alone 38.02 6.53 2.45 3.46
American Indian/Alaska Native alone 47.78 4.93 1.71 2.24
Asian alone 26.25 4.63 1.72 2.74
Native Hawaiian/OPI* alone 25.63 2.54 0.74 0.87
Some Other Race alone 20.16 5.36 2.28 3.21
Two or More Races 21.92 5.75 2.80 3.71
*OPI stands for Other Pacific Islander.
Table 10. Production US 2020 DHCH: Tenure by Householder Age $\rho$-zCDP MAE
Group States Counties Places Tracts
@@
Householder 15 to 24 years 112.76 13.67 4.52 3.50
Householder 25 to 34 years 130.73 22.13 8.09 8.31
Householder 35 to 54 years 142.14 27.59 10.05 10.86
Householder 55 to 64 years 130.59 24.99 8.65 9.77
Householder 65 years and over 151.25 23.70 9.14 10.55
@@
Householder 15 to 24 years 83.04 15.06 5.56 5.91
Householder 25 to 34 years 108.04 19.25 6.65 7.83
Householder 35 to 54 years 79.10 18.85 7.87 8.97
Householder 55 to 64 years 107.25 17.48 5.61 6.45
Householder 65 years and over 94.39 14.10 5.10 6.58
Table 11. Production US 2020 DHCH: Tenure by Race $\rho$-zCDP MAE
Group States Counties Places Tracts
Owner occupied
White alone 22.78 5.56 2.70 3.27
Black/African American alone 27.76 4.34 1.54 2.30
American Indian/Alaska Native alone 28.88 3.40 1.01 1.50
Asian alone 19.39 3.33 1.19 1.99
Native Hawaiian/OPI* alone 15.86 1.75 0.38 0.47
Some Other Race alone 13.33 3.81 1.60 2.24
Two or More Races 15.08 4.20 2.12 2.74
Renter occupied
White alone 17.08 4.70 2.26 2.95
Black/African American alone 13.92 3.94 1.42 2.26
American Indian/Alaska Native alone 20.80 2.83 0.99 1.47
Asian alone 12.10 2.67 0.95 1.63
Native Hawaiian/OPI* alone 13.47 1.59 0.42 0.61
Some Other Race alone 11.82 3.29 1.41 2.15
Two or More Races 12.24 3.77 1.66 2.49
@@
Table 12. Production US 2020 DHCH: Household Size $\rho$-zCDP MAE
Group States Counties Places Tracts
1-person household 18.10 18.83 8.07 7.92
2-person household 79.06 36.10 13.41 12.20
3-person household 65.67 29.59 10.55 11.67
4-person household 57.94 25.37 9.27 11.12
5-person household 51.45 24.91 9.10 9.35
6-person household 79.84 30.78 9.72 7.51
7-or-more-person household 98.96 31.56 9.55 6.89
Table 13. Production US 2020 DHCP: Relationship to Householder $\rho$-zCDP MAE
Group States Counties Places Tracts
Householder 23.02 6.70 4.32 4.26
Opposite-sex husband/wife/spouse 8.43 4.50 3.39 3.26
Opposite-sex unmarried partner 10.06 3.45 2.40 2.37
Same-sex husband/wife/spouse 8.39 2.58 1.44 1.47
Same-sex unmarried partner 6.92 2.33 1.37 1.35
Biological son or daughter 18.67 5.92 3.75 3.68
Adopted son or daughter 9.98 3.64 2.17 1.98
Stepson or stepdaughter 9.59 3.75 2.29 2.17
Brother or sister 8.35 3.30 2.15 2.18
Father or mother 7.02 3.08 2.08 2.06
Grandchild 11.35 3.91 2.60 2.58
Parent-in-law 6.25 2.63 1.47 1.57
Son-in-law or daughter-in-law 5.31 2.70 1.63 1.67
Other relative 11.02 3.93 2.33 2.33
Roommate or housemate 9.73 3.80 2.15 2.30
Foster child 6.61 2.58 1.46 1.23
Other nonrelative 10.80 3.84 2.37 2.21
Table 14. Production US 2020 DHCP: 5 Year Age Groups $\rho$-zCDP MAE
Group Counties Places Tracts AIR/ORT*
Under 5 years 9.52 4.47 5.28 3.95
5 to 9 years 10.64 4.36 5.84 4.35
10 to 14 years 10.63 4.50 5.95 4.61
15 to 19 years 8.31 4.52 4.39 2.84
20 to 24 years 4.57 2.78 3.06 2.08
25 to 29 years 4.59 2.80 3.09 2.17
30 to 34 years 4.39 2.77 3.02 2.12
35 to 39 years 6.56 3.01 4.06 3.29
40 to 44 years 6.36 2.94 4.00 2.86
45 to 49 years 6.25 2.93 3.95 2.89
50 to 54 years 6.22 2.99 3.96 3.20
55 to 59 years 6.17 3.07 3.96 3.35
60 to 64 years 8.08 3.80 4.99 3.33
65 to 69 years 7.90 3.63 4.75 3.53
70 to 74 years 6.01 2.84 3.67 2.83
75 to 79 years 5.87 2.71 3.50 2.76
80 to 84 years 5.88 2.55 3.33 2.54
85 to 89 years 5.39 2.29 3.10 1.90
90 to 94 years 5.29 2.19 2.75 1.55
95 to 99 years 4.77 1.69 2.00 0.99
100 to 104 years 3.14 0.94 0.77 0.38
105 to 109 years 1.29 0.27 0.14 0.40
over 109 years 0.64 0.03 0.05 0.19
@@
Tables 7 through 12 provide the $\rho$-zCDP MAE values of tabulations in the DHCH universe. As described previously, counts in this universe can be viewed as counts of households; alternatively, since there is one householder per household, counts in this universe can also be viewed as counts of householders. For the DHCP universe, Tables 13 and 14 provide the $\rho$-zCDP MAE of counts of individuals grouped by relationship to the householder and by their age group, respectively.
@@ 7. Conclusion and Progress after the 2020 Census DHC Release
As described in Section 2, one important advantage of formally private mechanisms like the DAS is that they are sufficiently transparent to allow users to construct confidence intervals of the confidential CEF-based tabulations in a straightforward manner. To describe how this can be done, more background on how the strategy matrices described in Section 5 are used within the DAS will be helpful. Specifically, after spine optimization, the noisy measurements for each tabulation and each geographic unit are defined by adding discrete Gaussian noise to the CEF-based tabulation of the given geographic unit. If the PLB allocation of the geographic unit and tabulation pair is given by $\rho,$ the mean-zero noise added to each level of the CEF-based tabulation is given by a realization of a random variable from the discrete Gaussian distribution, $N_{\mathbb{Z}}(\mu=0,\sigma^{2}=1/\rho),$ which is defined so that its probability mass function is proportional to $\exp(-x^{2}/2\sigma^{2})$ for each $x\in\mathbb{Z}$ and zero otherwise [canonne2020discrete]. We call the collection of all such noisy measurements for a given DAS execution the noisy measurement files (NMFs).
To support research and development on tools that would provide users inferences on underlying CEF-based tabulations, the Census Bureau has released NMFs for both the final 2010-CEF-based DDP production settings DAS implementations and the 2020 production DAS implementations as experimental data products. Specifically, Table 15 provides AWS Open Data URLs for the DAS evaluation NMFs using our 2020 production settings, for both the Redistricting Data File and the DHC statistical data products and for both the 2020 production DAS executions and the DDPs DAS executions that used 2010 input data. Each of these four sites contains a README file that provides information on the format of the corresponding NMFs, an AWS Simple Storage Service (S3) path for the NMFs, and instructions on how to download these NMFs using the AWS S3 command line interface.
Table 15. URLs for DAS NMFs hosted on AWS Open Data, for both the 2010 and $20$ input data vintages and for both the DHC and redistricting data file statistical data products
Data Product Vintage URL
Redistricting 2010 https://registry.opendata.aws/census-2010-pl94-nmf/
2020 https://registry.opendata.aws/census-2020-pl94-nmf/
DHC 2010 https://registry.opendata.aws/census-2010-dhc-nmf/
2020 https://registry.opendata.aws/census-2020-dhc-nmf/
Subsets of the NMFs that provide query answers for queries that are at least as granular as the query of interest can be used to generate point estimates and confidence intervals of the confidential CEF-based tabulations via either simulation or a continuous Gaussian modeling assumption; see for example [nmfCIs, McCartan2023Making]. Alternatively, rather than using small subsets of the noisy measurements from the NMFs, [cumings2024full] provides a full-information confidence interval estimator, i.e., an estimator that uses all noisy measurements from the respective DAS execution. The Census Bureau is currently exploring releasing confidence intervals of CEF-based tabulations that are computed using an implementation of this approach as an experimental data product.
Confidence interval estimators based on the NMFs, such as the examples cited above, result in intervals that are not centered around the count implied by the MDF output from the DAS, and instead are centered around an unbiased estimate of the CEF-based tabulation. In contrast, [ashmead2024] provide several heuristic confidence interval estimators centered directly around the published tabulations that are based on MDFs generated by simulating the DHC DAS implementations, with the input of each simulation iterate being either the 2020 production DHC DAS output MDFs or DHC DAS output MDFs that used a 2010-vintage input CEF. By using MDFs that have already been protected by a formally private mechanism as input rather than the confidential CEF, this method avoids additional PLB expenditures. In some cases these confidence interval estimates are more narrow than those that are based on creating an unbiased estimator of the target query using the 2020 production DHC DAS NMFs. The Census Bureau has released the simulation MDFs required to compute these confidence interval estimates, and [ashmead2024] provide instructions for downloading these simulation MDFs.